Vitamin C Project Lab: Part B: Data analysis Calculate the grams of Vitamin C in your Erlenmeyer. (as in titration #1 ) Calculate the grams of Vitamin C per tablet. Calculate a percent error by comparing your result to the tablet dose above. Determine the % composition of vitamin C in the tablet Build a data table similar to the one you used for titration # 1 to calculate the following: 1. 2 3. 4. Show all your work Trial Triel1 Grams af 0.859 6580 Vitamin Initial vollme f NoOH Hindl Voleme of Na Oh valeme 32. 70 mL 32.50 ML
Part B: Titration Vitamin C tablet-Background Vitamin C, ascorbic acid (CeHsO6), is commonly sold in tablet form with a starch binder Thus, vitamin C tablets are mixtures. Tablets are labeled according to their vitamin C content, not according to their weight. Therefore, a 1 gram tablet would have less than 1 gram of Vitamin C. You will chemically analyze vitamin C tablets to determine their actual vitamin C content Procedure: You will keep records of this experiment in your lab notebook. Follow the notebook directions given to you by your instructor. 1. Obtain from your teacher the brand name of Vitamin C tablets you'll be using and the tablet dose. Calculate the number of tablets you will need for 500 mg of Vitamin C. Complete this table in your lab notebook. Brand name of tablet: now Tablet Dose (mg vitamin C per tablet): Number of tablets used to get 500-600 mg sample: 2 Read through the directions (steps 3-10) and create a data table in your lab notebook in which you can collect your data. 3. Weigh out a sample of one brand of whole vitamin C tablets so that the sample contains around 500 mg of vitamin C (1, 2, or as many tablets as are needed). The tablets usually weigh more than the dose size because of binders added to help hold the tablet together. 4. Use a mortar and pestle to grind the tablets into a powder. Transfer each sample into a rinsed and marked 250 ml flask 6 5. Add about 50 ml deionized water and 2 drops of phenolphthalein indicator to each sample. 7 Add a magnetic stir bar into your sample and use a heating plate with magnetic spinner to stir your sample (do not heat the sample), you may use this stirring it while proceeding with the titration. 8 Titrate your sample with 0.100M NaOH, following the directions given in the previous section. (Dyes in the tablet may cause difficulty in observing the endpoint.) 9. Retrieve the magnet with a magnetic stick and place it in the next sample. Return it to the cart once you finish. 10. Record the concentration of NaOH, the initial volume of NAOH in the buret and the final volume of NaOH in the buret. 11. Run two trials in total.
Answer #1

Part B

1) Grams of Vitamin C in Erlenmeyer

Volume of NaOH in Erlenmeyer

Trail 1: Final - Initial = 32.50 - 0 = 32.50 ml

Trail 2: Final - Initial = 32.70 - 0 = 32.70 ml

Moles of NaOH in Erlenmeyer (molarity is 0.100M)

Trail 1: moles of NaOH = 0.100 M × 0.0325 L = 0.00325 mol

Trail 2: moles of NaOH = 0.100 M × 0.0327 L = 0.00327 mol

As reaction ratio of NaOH and Vitamin C is 1:1, so NaOH reacted with same moles of Vitamic C, as moles of NaOH

Trail 1: moles of Vitamin C = 0.100 M × 0.0325 L = 0.00325 mol

Trail 2: moles of Vitamin C = 0.100 M × 0.0327 L = 0.00327 mol

Since molar mass of Vitamin C is 176.12 g/mol

grams = moles × molar mass

Trail 1: grams of Vitamin C = 176.12 g/mol × 0.00325 mol = 0.57 g

Trail 2: grams of Vitamin C = 176.12 g/mol × 0.00327 mol = 0.58 g

2) Grams of Vitamin C per tablet

500mg of 1 tablet Vitamin C contains less than 1 g of Vitamin C. From above calculation

Trial 1: grams of Vitamin C per tablet is 0.57 g

Trial 2: grams of Vitamin C per tablet is 0.58 g

3) %Error of grams of Vitamin C per tablet to the tablet dose

%error = (grams of Vitamin C - tablet dose ) / tablet dose × 100

Trail 1: %error = (0.57 g - 0.5 g) / 0.5 g × 100 = 14%

Trail 2: %error = (0.58 g - 0.5 g) / 0.5 g × 100 = 16%

4) %composition of Vitamin C per tablet is calculated using below equation

%composition of Vitamin C = (mass of Vitamin C per tablet / mass of 1 tablet given ) × 100

Trial 1: %composition of Vitamin C = (0.57g / 0.859)  × 100 = 66.36 %

Trial 2: %composition of Vitamin C = (0.58g / 0.859)  × 100 = 67.52 %